OpenStudy (anonymous):

\[A=\left[\begin{matrix}1 & 0 \\ 3& 2\end{matrix}\right], B=\left[\begin{matrix}x & 0 \\ 1 & 3\end{matrix}\right]\] and AB=BA find x.

OpenStudy (phi):

Can you multiply out A*B?

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get?

OpenStudy (anonymous):

\[\left[\begin{matrix}x & 0 \\ 3x+2 & 6\end{matrix}\right]\]

OpenStudy (phi):

most excellent. now do B*A

OpenStudy (anonymous):

\[\left[\begin{matrix}x & 0 \\ 10 & 6\end{matrix}\right]\]

OpenStudy (phi):

now the problem says those two matrices are equal (corresponding elements are equal) x=x yes 0= 0 yes 6=6 yes 3x+2=10 this one is interesting. solve for x

OpenStudy (anonymous):

i see.

OpenStudy (anonymous):

uber cool